OpenStudy (anonymous):

Given the functions f(x) = 2x-7/x+1 and g(x) =2√x find: A) (f º g)(x) and state domain B) (g º f)(x) and state domain C) (f º g) (9)

OpenStudy (ghazi):

to find (fog)(x)= just substitute the value of g(x) at x in f(x) and in (gof) use value of f(x) in g(x)...

OpenStudy (anonymous):

As @ghazi says: \[(f \ o\ g)(x) = \frac{4\sqrt{x}-7}{\sqrt{x}+1}\] \[(g\ o\ f)(x) = 2\sqrt {\frac{2x -7}{x+1}}\] Domain for the first is not restricted since sqrt(x) >= 0 for all x. Domain for the second: x cannot be equal to -1. Compute (f o g)(9) = (12-7)/(3 + 1) = 5/4

OpenStudy (anonymous):

Oh, and additional restriction for the second (2x - 7) >= 0 only.

OpenStudy (anonymous):

Oh, and additional restriction for the first: x >= 0.